Ray GI bisects angle DGH so that m angle DGI is x – 3 and m angle IGH is 2x – 13. Find the value of x.
Answer by cleomenius(959)

The ray bisects the angle DGH so DGI will equal IGH.
x - 3 = 2x - 13
10 = x
DGI = 10 - 3 = 7.
IGH = 20 - 13 = 7.
Cleomenius.